OpenStudy (anonymous):

If f(2)=5 and f(-1)=-3, find a and b in function f(x)=ax+b.

OpenStudy (mertsj):

You are given two points on the line and asked to find the equation of the line. So first find the slope.

OpenStudy (anonymous):

If im to find the slope ... then what would be the coordinates. Unless the coordinates are (2,5) and (-1,-3) Thats telling me f(x)=y ......

OpenStudy (mertsj):

Yes. Those are the ordered pairs. Find the slope.

OpenStudy (anonymous):

f(2) = 2a+b=5 f(-1)=-a+b=-3

OpenStudy (anonymous):

then b+3 =a

OpenStudy (anonymous):

@tianz2303 thats what I figure the answer would be !!!!! ..... but I have doubts ! :(

OpenStudy (anonymous):

then you keep going like 2(b+3) + b=5

OpenStudy (anonymous):

then b= -(1/3)

OpenStudy (anonymous):

if a = b +3 then a= -(1/3) + 3 which is ?

OpenStudy (anonymous):

a=8/3 so there you go :D

OpenStudy (anonymous):

is it good?

OpenStudy (anonymous):

Yeah! It seems good. Just ... a bit confusing. Im trying to break it down.

OpenStudy (anonymous):

lol lemme put everything together

OpenStudy (anonymous):

f(2) = 2a+b=5 f(-1)=-a+b=-3 then b+3 =a which would help us replace into first equation right? 2(b+3) + b=5 if you solve it 2b + 6 + b = 5 which will make 3b = -1 and finally b = - 1/3 if we go back to b+3 = a we replace our b value and we have a = 8/3 finally a=8/3 b= - 1/3

OpenStudy (anonymous):

Totally ! Totally ! Tooooootaaaaally !!!!!! I totally got it ! I did it and it came out easy !!! It's like finding the equation of a line just .... differently !!!

OpenStudy (anonymous):

TOTALLY !!! Your awesome!

OpenStudy (anonymous):

:D yaaay glad i helped!

OpenStudy (mertsj):

Two points: (2,5) and (-1,-3) Slope = \[m=\frac{5-(-3)}{2-(-1)}=\frac{8}{3}\] Point slope form: \[y-y _{1}=m(x-x _{1})\] Plug in the slope and either point: \[y-5=\frac{8}{3}(x-2)\] \[y=\frac{8}{3}x-\frac{1}{3}\]

OpenStudy (anonymous):

There we go @Mertsj , totally on point ! I did it this way just seconds ago and my answers matched. I find that doing it @Mertsj way will be easier for me on my test! Awesome. Thanks guys.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

haha sure i just go the easy way :/